slove the system by addition
x+y=8,x-y=8
i come up with 0
can someone explained this to me please

Answer by scott8148(6628)

adding the equations eliminates the y term ___ (x + y) + (x - y) = 8 + 8 ___ 2x = 16 ___ x = 8

substituting for x ___ (8) + y = 8 ___ y = 0 (as you already know)